### v36

# A Simple Proof of the Riemann Hypothesis

Abstract
In this article, it is proved that the non-trivial zeros of the Riemann zeta function must lie on the critical line, known as the Riemann hypothesis.
Keywords Riemann zeta function $\cdot$ Riemann hypothesis  $\cdot$ Non-trivial zeros  $\cdot$ Critical line
1 Riemann Zeta function
The Riemann zeta function is defined over the complex plane as [1],
$\zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n^{s}}\quad,\Re(s)>1$ (1)
where $\Re(s)$ denotes the real part of $s$. There are several forms can be used for an analytic continuation for $\Re(s)>0$ such as [1, 2],
$\zeta(s)=\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{N^{1-s}}{1-s}-s\int_{N}^{\infty}\frac{x-\lfloor x\rfloor}{x^{s+1}}dx,N\in\mathbb{N}$ (2)
where $\lfloor x\rfloor$ is the floor or integer part such that $x-1<\lfloor x\rfloor\leq x$ for real $x$.
and
$\zeta(s)=\frac{1}{\left(1-2^{1-s}\right)}\eta(s),\qquad\eta(s)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n^{s}}$ (3)
where $s\neq 1+\frac{2\pi ki}{\log(2)},k=0,\pm 1,\pm 2\ldots$ and $\eta(s)$ is the Dirichlet eta function (sometimes called the alternating zeta function).
Using Euler-Maclaurin formula, it can also be written as [1, 2],
$\zeta(s)=\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\sum_{i=1}^{m}\binom{s+2i-2}{2i-1}\frac{B_{2i}}{2i}N^{1-s-2i}-\epsilon_{m},\quad\Re(s)>-2m;m,N\in\mathbb{N}$ (4)
where the error term is given by,
$\epsilon_{m}$ $=\binom{s+2m}{2m+1}\int_{N}^{\infty}\frac{\bar{B}_{2m+1}(x)}{x^{s+2m+1}}dx,$ (5)
$B_{k}$ is the $k$-th Bernoulli number defined implicitly by,
$\frac{t}{e^{t}-1}=\sum_{k=0}^{\infty}B_{k}\frac{t^{k}}{k!},$ (6)
$B_{k}(x)$ is the $k$-th Bernoulli polynomial defined as the unique polynomial of degree $k$ with the property that,
$\int_{t}^{t+1}B_{k}(x)dx=t^{k},$ (7)
$\bar{B}_{k}(x)$ is the periodic function $B_{k}(x-\lfloor x\rfloor)$.
Riemann also extended $\zeta(s)$ to $\mathbb{C}$ as a meromorphic function, with only a simple pole at $s=1$ with residue 11, by the functional equation,
$\zeta({s})=2^{s}\pi^{s-1}\sin\left(\frac{\pi s}{2}\right)\Gamma\left({1-s}\right)\zeta({1-s}),\qquad s\in\mathbb{C}\setminus\{0,1\}$ (8)
2 Zeros of the Riemann Zeta Function
The trivial zeros of the Riemann zeta function occur at the negative even integers; that is, $\zeta(-2n)=0,n\in\mathbb{N}$ [1]. On the other hand, the non-trivial zeros lie in the critical strip, $0\leq\Re(s)\leq 1$. Both Hadamard [3] and de la Vallee Poussin [4] independently proved that there are no zeros on the boundaries of the critical strip (i.e. $\Re(s)=0$ or $\Re(s)=1$). Gourdon and Demichel [5] verified the Riemann Hypothesis until the $10^{13}$-th zero.
Mossinghoff and Trudgian [6] proved that there are no zeros for $\zeta(\sigma+it)$ for $|t|\geq 2$ in the region,
$\sigma\geq 1-\frac{1}{5.573412\log|t|}$ (9)
This represents the largest known zero-free region for the zeta-function within the critical strip for $3.06\times 10^{10}<|t|<\exp(10151.5)\approx 5.5\times 10^{4408}$.
Due to the functional equation (8) and the complex conjugation properties, the non-trivial zeros of $\zeta(s)$ are symmetric with respect to both the critical line and the real axis, that is $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$. According to equation (3), the zeros of the Dirichlet eta function include all the non-trivial zeros of the zeta function. So, if $s$ is a non-trivial zero of the zeta function, then $\eta(s)=\eta(1-s)=\eta(\bar{s})=\eta(1-\bar{s})=0$.
3 Riemann Zeta Function and Euler-Maclaurin Formula
Prior to proving the Riemann hypothesis, let us first prove the following theorems inspired by the Euler-Maclaurin formula for the zeta function.
Theorem 1.
Assume that $s=\sigma+it$ for $0<\sigma<1,t\in\mathbb{R}$, $N\in\mathbb{N}$ and
$\zeta_{N}(s):=\sum_{n=1}^{N}\frac{1}{n^{s}},$ (10)
$\phi_{N}(s):=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}$ (11)
Then:
  1. i)
    $N^{s}\zeta(s)=N^{s}\phi_{N}(s)-\frac{1}{2}+\frac{s}{12N}+O\left(\frac{1}{N^{2}}\right)$,
  2. ii)
    $\zeta(s)=0\Rightarrow\lim_{N\to\infty}N^{s}\phi_{N}(s)=1/2$,
  3. iii)
    $\zeta(s)\neq 0\Rightarrow N^{s}\phi_{N}(s)=O\left(N^{\sigma}\right)\textnormal{~as~}N\to\infty$.
Proof.
  1. i)
    Using $m=1$ in equations (4) and (5), we get,
    $\zeta(s)$ $=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}-\epsilon_{1}=\phi_{N}(s)-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}-\epsilon_{1}$ (12)
    and the error term is given by,
    $\epsilon_{1}$ $=\binom{s+2}{3}\int_{N}^{\infty}\frac{\bar{B}_{3}(x)}{x^{s+3}}dx,$ (13)
    Since,
[…]
    then, the error term, $\epsilon_{1}$, can be bounded as,
    $|\epsilon_{1}|$ $\leq B^{max}_{3}\frac{|s(s+1)(s+2)|}{3!}\int_{N}^{\infty}\frac{1}{x^{\sigma+3}}dx\leq B^{max}_{3}\frac{|s(s+1)(s+2)|}{(3!)(\sigma+2)}\left(\frac{1}{N^{\sigma+2}}\right)$ (15)
    Therefore,
    $\zeta(s)=\phi_{N}(s)-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}+O\left(\frac{1}{N^{{}^{\sigma+2}}}\right)$ (16)
    Multiplying by $N^{s}$ leads to,
    $N^{s}\zeta(s)=N^{s}\phi_{N}(s)-\frac{1}{2}+\frac{s}{12N}+O\left(\frac{1}{N^{{}^{2}}}\right)$ (17)
  2. ii)
    At $\zeta(s)=0$,
    $0=N^{s}\phi_{N}(s)-\frac{1}{2}+\frac{s}{12N}+O\left(\frac{1}{N^{{}^{2}}}\right)$ (18)
    Hence,
    $\lim_{N\to\infty}N^{s}\phi_{N}(s)=\frac{1}{2}$ (19)
    Note that this property can be numerically verified at the well-known non-trivial zeros of zeta function using any software package. For example, the following MATLAB [7] code verifies this property for the first two non-trivial zeros of $\zeta(s)$, which are $s1\approx 0.5+14.1347251i$ and $s2\approx 0.5+21.0220396i$,
    >>>> $s1=0.5+14.1347251i;N=1e6;n=1:N;\text{sum}((N./n).\text{\^{}}s1)-N/(1-s1)$
    >>>> $s2=0.5+21.0220396i;N=1e6;n=1:N;\text{sum}((N./n).\text{\^{}}s2)-N/(1-s2)$
    which produce, 0.5000.000i0.500-0.000i.
  3. iii)
    At $\zeta(s)\neq 0$, as $N\to\infty$,
    $N^{s}\zeta(s)=O\left(N^{\sigma}\right)$ (20)
    and according to equation (17),
    $N^{s}\phi_{N}(s)=N^{s}\zeta(s)+\frac{1}{2}+O\left(\frac{1}{N}\right)=O\left(N^{\sigma}\right)$ (21)
Theorem 2.
Under the same assumptions and definitions of Theorem˜1, suppose further that,
$\eta_{N}(s):=\sum_{n=1}^{N}\frac{(-1)^{n+1}}{n^{s}},$ (22)
$\nu_{N}(s):=N^{s}\eta_{N}(s)-N^{1-\bar{s}}\eta_{N}(1-\bar{s}),$ (23)
$\mu_{N}(s):=N\nu_{N}(s)$ (24)
Then, as $N\to\infty$,
  1. i)
    $N^{s}\eta_{N}(s)\to\begin{cases}-1/2&\text{if }N\text{ is even}\\ 1/2&\text{if }N\text{ is odd}\end{cases}$,
  2. ii)
    $\zeta(s)\neq 0\Rightarrow N^{s}\eta_{N}(s)=O(N^{\sigma})$,
  3. iii)
    $\sigma=1/2\Rightarrow\nu_{N}(s)=\mu_{N}(s)=0$,
  4. iv)
    $\zeta(s)\neq 0\textnormal{~and~}\sigma\neq 1/2\Rightarrow\nu_{N}(s),\mu_{N}(s)\textnormal{~diverge}$,
  5. v)
    $\zeta(s)=0\Rightarrow\nu_{N}(s)\to 0,\mu_{N}(s)\to\begin{cases}(2\sigma-1)/4&\text{if }N\text{ is even}\\ (1-2\sigma)/4&\text{if }N\text{ is odd}\end{cases}$.
Proof.
  1. i)
    Assume that NN is even and let us express $\eta_{N}(s)$ in terms of $\zeta_{N}(s)$ as follows,
    $\eta_{N}(s)=\sum_{n=1}^{N}\frac{(-1)^{n+1}}{n^{s}}=\sum_{n=1}^{N}\frac{1}{n^{s}}-2\sum_{n=1}^{N/2}\frac{1}{(2n)^{s}}=\zeta_{N}(s)-2^{1-s}\zeta_{N/2}(s)$ (25)
    Multiplying by $N^{s}$,
    $N^{s}\eta_{N}(s)=N^{s}\zeta_{N}(s)-2^{1-s}N^{s}\zeta_{N/2}(s)$ (26)
    Using equation (11),
    $N^{s}\zeta_{N}(s)=N^{s}\left[\phi_{N}(s)+\frac{N^{1-s}}{1-s}\right]=N^{s}\phi_{N}(s)+\frac{N}{1-s}$ (27)
    Similarly,
    $N^{s}\zeta_{N/2}(s)=N^{s}\left[\phi_{N/2}(s)+\frac{(N/2)^{1-s}}{1-s}\right]=N^{s}\phi_{N/2}(s)+\frac{N}{2^{1-s}(1-s)}$ (28)
    Substituting in equation (26),
    $N^{s}\eta_{N}(s)=N^{s}\phi_{N}(s)+\frac{N}{1-s}-2^{1-s}\left[N^{s}\phi_{N/2}(s)+\frac{N}{2^{1-s}(1-s)}\right]=N^{s}\phi_{N}(s)-2^{1-s}N^{s}\phi_{N/2}(s)$ (29)
    At $\zeta(s)=0$, from equation (19),
    $\lim_{N\to\infty}N^{s}\phi_{N/2}(s)=2^{s}\lim_{N\to\infty}\left(\frac{N}{2}\right)^{s}\phi_{N/2}(s)=2^{s}\left(\frac{1}{2}\right)=\frac{1}{2^{1-s}}$ (30)
    Thus, substituting by equations (19) and (30) in equation (29) as $N\to\infty$ implies that,
    $N^{s}\eta_{N}(s)\to\frac{1}{2}-2^{1-s}\left(\frac{1}{2^{1-s}}\right)=\frac{1}{2}-1=-\frac{1}{2}$ (31)
    For odd NN,
    $N^{s}\eta_{N}(s)$ $=N^{s}\sum_{n=1}^{N}\frac{(-1)^{n+1}}{n^{s}}=N^{s}\left[\sum_{n=1}^{N-1}\frac{(-1)^{n+1}}{n^{s}}+\frac{1}{N^{s}}\right]$ (32)
    $=\left[\frac{N^{s}}{(N-1)^{s}}\right]\left[(N-1)^{s}\sum_{n=1}^{N-1}\frac{(-1)^{n+1}}{n^{s}}\right]+1\to\left(1\right)\left(-\frac{1}{2}\right)+1=\frac{1}{2}$
    Thus, equations (31) and (32) can be combined for $N\to\infty$ as
[…]
    This can also be numerically verified using the following MATLAB code for the first two non-trivial zeros of $\zeta(s)$ for even NN,
    >>>> $s1=0.5+14.1347251i;N=1e6;n=1:N;\text{sum}((-1).\text{\^{}}(n+1).*(N./n).\text{\^{}}s1)$
    >>>> $s2=0.5+21.0220396i;N=1e6;n=1:N;\text{sum}((-1).\text{\^{}}(n+1).*(N./n).\text{\^{}}s2)$
    which produce, 0.5000.000i-0.500-0.000i
    and for odd NN,
    >>>> $s1=0.5+14.1347251i;N=1e6+1;n=1:N;\text{sum}((-1).\text{\^{}}(n+1).*(N./n).\text{\^{}}s1)$
    >>>> $s2=0.5+21.0220396i;N=1e6+1;n=1:N;\text{sum}((-1).\text{\^{}}(n+1).*(N./n).\text{\^{}}s2)$
    which produce, 0.5000.000i0.500-0.000i
    (a) $t=14.1347251.$ $\sigma=0.7$ ( ), $\sigma=0.3$ ( ).
    (b) $t=49.7738325.$ $\sigma=0.7$ ( ), $\sigma=0.3$ ( ).
    (c) $t=14.1347251.$ $\sigma=0.6$ ( ), $\sigma=0.4$ ( ).
    (d) $t=49.7738325.$ $\sigma=0.6$ ( ), $\sigma=0.4$ ( ).
    (e) $s=0.5+14.1347251i$.
    (f) $s=0.5+49.7738325i$.
    Figure 1: Graphs of $\Re\{N^{s}\eta_{N}(s)\}$ where $s=\sigma+it$ for different values of NN ranges from 11 to 200200.
  2. ii)
    Suppose that NN is even, then from equation (16),
    $\zeta(s)=\phi_{N/2}(s)-\frac{2^{s}}{2N^{s}}+\frac{2^{s+1}s}{12N^{s+1}}+O\left(\frac{1}{N^{{}^{\sigma+2}}}\right)$ (34)
    Multiplying by $N^{s}$,
    $N^{s}\zeta(s)=N^{s}\phi_{N/2}(s)-\frac{1}{2^{1-s}}+\frac{2^{s+1}s}{12N}+O\left(\frac{1}{N^{{}^{2}}}\right)$ (35)
    Substituting by equations (17) and (35) in equation (29),
    $N^{s}\eta_{N}(s)$ $=N^{s}\zeta(s)+\frac{1}{2}-\frac{s}{12N}+O\left(\frac{1}{N^{{}^{2}}}\right)-2^{1-s}\left[N^{s}\zeta(s)+\frac{1}{2^{1-s}}-\frac{2^{s+1}s}{12N}+O\left(\frac{1}{N^{{}^{2}}}\right)\right]$ (36)
    $=-\frac{1}{2}+N^{s}\zeta(s)\left[1-2^{1-s}\right]+\frac{s}{4N}+O\left(\frac{1}{N^{{}^{2}}}\right)$
    $=-\frac{1}{2}+N^{s}\eta(s)+\frac{s}{4N}+O\left(\frac{1}{N^{{}^{2}}}\right)$
    Thus, at $\zeta(s)\neq 0$, as $N\to\infty$,
    $N^{s}\eta_{N}(s)=-\frac{1}{2}+N^{s}\eta(s)=O(N^{\sigma})$ (37)
    For odd NN, using equation (36)
    $N^{s}\eta_{N}(s)$ $=N^{s}\left[\sum_{n=1}^{N-1}\frac{(-1)^{n+1}}{n^{s}}+\frac{1}{N^{s}}\right]=\frac{N^{s}}{(N-1)^{s}}\left[(N-1)^{s}\sum_{n=1}^{N-1}\frac{(-1)^{n+1}}{n^{s}}\right]+1$ (38)
    $=\left(\frac{N}{N-1}\right)^{s}\left[-\frac{1}{2}+(N-1)^{s}\eta(s)+\frac{s}{4(N-1)}+O\left(\frac{1}{(N-1)^{{}^{2}}}\right)\right]+1$
    $\to\frac{1}{2}+(N-1)^{s}\eta(s)=O(N^{\sigma})$
    These properties can be visualized in Figure 1 which shows $\Re\{N^{s}\eta_{N}(s)\}$ for different values of NN ranges from 11 to 200200 for $s=\sigma+14.1347251i$ in Figure 1(a), Figure 1(c), Figure 1(e) and for $s=\sigma+49.7738325i$ in Figure 1(b), Figure 1(d), Figure 1(f). It can be noticed that the sum oscillates as NN increases without bound for $\sigma\neq 1/2$. However, at $\zeta(s)=0$ (Figure 1(e), Figure 1(f)), the sum is bounded, and as NN increases, it tends to alternate between 1/2-1/2 for even NN and $1/2$ for odd NN.
  3. iii)
    At $\sigma=1/2$, we have $s=1-\bar{s}$, hence
    $\nu_{N}(s)=N^{s}\eta_{N}(s)-N^{1-\bar{s}}\eta_{N}(1-\bar{s})=N^{s}\eta_{N}(s)-N^{s}\eta_{N}(s)=0$ (39)
    and similarly,
    $\mu_{N}(s)=N\nu_{N}(s)=0$ (40)
    That is, $\nu_{N}(s)$ and $\mu_{N}(s)$ vanish at all points along the critical line, $\sigma=1/2$, including the well-known non-trivial zeros, $\zeta(s)=0$, and the non-zeros, $\zeta(s)\neq 0$.
  4. iv)
    At $\zeta(s)\neq 0\textnormal{~and~}\sigma\neq 1/2$,
    From equation (37), we have,
    $\nu_{N}(s)=N^{s}\eta_{N}(s)-N^{1-\bar{s}}\eta_{N}(1-\bar{s})=O\left(N^{\sigma}\right)-O\left(N^{1-\sigma}\right)$ (41)
    which diverges as $N\to\infty$ since either $N^{\sigma}$ or $N^{1-\sigma}$ will dominate.
    Consequently,
    $\mu_{N}(s)=N\nu_{N}(s)=O\left(N^{1+\sigma}\right)-O\left(N^{2-\sigma}\right)$ (42)
    which diverges too.
  5. v)
    At $\zeta(s)=0$,
    Due to the symmetry of the non-trivial zeros of the zeta function about the critical line, at $\zeta(1-\bar{s})=0$, and according to equation (33), as $N\to\infty$, we must also have,
[…]
    Thus, as $N\to\infty$,
    $\nu_{N}(s)=\left[N^{s}\eta_{N}(s)-N^{1-\bar{s}}\eta_{N}(1-\bar{s})\right]\to 0$ (44)
    For even NN, using equation (36),
    $\mu_{N}(s)$ $=N\left[-\frac{1}{2}+N^{s}\eta(s)+\frac{s}{4N}+O\left(\frac{1}{N^{{}^{2}}}\right)\right]-N\left[-\frac{1}{2}+N^{1-\bar{s}}\eta(1-\bar{s})+\frac{1-\bar{s}}{4N}+O\left(\frac{1}{N^{{}^{2}}}\right)\right]$ (45)
    $\to\frac{2\sigma-1}{4}\text{ as }N\to\infty$
    Similarly, for odd NN, using equation (38),
    $\mu_{N}(s)$ $=N\left\{\left(\frac{N}{N-1}\right)^{s}\left[-\frac{1}{2}+(N-1)^{s}\eta(s)+\frac{s}{4(N-1)}+O\left(\frac{1}{(N-1)^{{}^{2}}}\right)\right]+1\right\}$ (46)
    $\quad-N\left\{\left(\frac{N}{N-1}\right)^{1-\bar{s}}\left[-\frac{1}{2}+(N-1)^{1-\bar{s}}\eta(1-\bar{s})+\frac{1-\bar{s}}{4(N-1)}+O\left(\frac{1}{(N-1)^{{}^{2}}}\right)\right]+1\right\}$
    $=N\left\{\left[1+\frac{s}{N}+O\left(\frac{1}{N^{2}}\right)\right]\left[-\frac{1}{2}+(N-1)^{s}\eta(s)+\frac{s}{4(N-1)}+O\left(\frac{1}{(N-1)^{{}^{2}}}\right)\right]+1\right\}$
    $\quad-N\left\{\left[1+\frac{{1-\bar{s}}}{N}+O\left(\frac{1}{N^{2}}\right)\right]\left[-\frac{1}{2}+(N-1)^{1-\bar{s}}\eta(1-\bar{s})+\frac{1-\bar{s}}{4(N-1)}+O\left(\frac{1}{(N-1)^{{}^{2}}}\right)\right]+1\right\}$
    $\to\frac{1-2\sigma}{4}\text{ as }N\to\infty$
    Therefore, as $N\to\infty$,
[…]
4 Riemann Hypothesis
All the non-trivial zeros of the Riemann zeta function lie on the critical line $\Re(s)=1/2$.
Proof.
Assume that $\zeta(s)=0$ where $0<\sigma<1,t\in\mathbb{R}$ and all the definitions of Theorem˜1 and Theorem˜2 hold.
Without loss of generality, assume that NN is even. Using Perron’s formula [8],
$\eta_{N}(s)$ $=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{N^{z}}{z}\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n^{s+z}}dz+\frac{(-1)^{N+1}}{2N^{s}}=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{N^{z}\eta(s+z)}{z}dz-\frac{1}{2N^{s}},\qquad c>1-\sigma$ (48)
Using $c=1$, multiplying by $N^{s}$, and taking the limit as $N\to\infty$,
$\lim_{N\to\infty}N^{s}\eta_{N}(s)=\lim_{N\to\infty}\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\eta(s+z)}{z}dz-\frac{1}{2}$ (49)
According to equation (31), we must have,
$\lim_{N\to\infty}\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\eta(s+z)}{z}dz=\lim_{N\to\infty}\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{s+1+iy}\eta(s+1+iy)}{1+iy}dy=0$ (50)
Let
$f_{N}(s,z)=\frac{N^{s+z}\eta(s+z)}{z}$ (51)
Using analytical continuation, $f_{N}(s,z)$ can be written as,
$f_{N}(s,z)=\frac{N^{s+z}\left(1-2^{1-s-z}\right)\zeta(s+z)}{z}$ (52)
which has a simple pole at $z=0$. Let us take a rectangular contour $C$ as,
  1. i)
    The vertical line from 1iT1-iT to $1+iT$,
  2. ii)
    The horizontal line from $1+iT$ to $\sigma+iT$,
  3. iii)
    The vertical line from $\sigma+iT$ to $\sigma-iT$,
  4. iv)
    The horizontal line from $\sigma-iT$ to 1iT1-iT.
where $T\to\infty$.
So, by applying Cauchy residue theorem [9],
$\ointctrclockwise_{C}f_{N}(s,z)\,dz$ $=\int_{1-iT}^{1+iT}f_{N}(s,z)\,dz+\int_{1+iT}^{-\sigma+iT}f_{N}(s,z)\,dz+\int_{-\sigma+iT}^{-\sigma-iT}f_{N}(s,z)\,dz+\int_{-\sigma-iT}^{1-iT}f_{N}(s,z)\,dz$ (53)
$=2\pi i\operatorname{Res}_{z=0}f_{N}(s,z)=N^{s}\eta(s)=0$
The second integral can be written as,
$\int_{1+iT}^{-\sigma+iT}f_{N}(s,z)\,dz=\int_{1}^{-\sigma}\frac{N^{s+x+iT}\left(1-2^{1-s-x-iT}\right)\zeta(s+x+iT)}{x+iT}\,dx$ (54)
and we have,
$|x+iT|\geq T,\qquad\left|1-2^{1-s-x-iT}\right|\leq 3$ (55)
From the convexity property of the zeta function [10], for large $T$, we have
[…]
Therefore,
$\int_{1+iT}^{-\sigma+iT}f_{N}(s,z)\,dz\to 0\text{ as }T\to\infty$ (57)
and similarly, the fourth integral,
$\int_{-\sigma-iT}^{1-iT}f_{N}(s,z)\,dz\to 0\text{ as }T\to\infty$ (58)
Substituting in equation (53),
$\int_{1-i\infty}^{1+i\infty}f_{N}(s,z)\,dz$ $=\int_{-\sigma-i\infty}^{-\sigma+i\infty}f_{N}(s,z)\,dz=0$ (59)
$\Rightarrow\lim_{N\to\infty}\int_{-\infty}^{\infty}\frac{N^{i(t+y)}\eta(i(t+y))}{-\sigma+iy}dy=0$
Due to the symmetry of the non-trivial zeros of the zeta function about the critical line, $\zeta(1-\bar{s})=0$ and similarly, we must have,
$\lim_{N\to\infty}N^{1-\bar{s}}\eta_{N}(1-\bar{s})=\lim_{N\to\infty}\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{1-\bar{s}+z}\eta(1-\bar{s}+z)}{z}dz-\frac{1}{2}$ (60)
and
$\lim_{N\to\infty}\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{1-\bar{s}+z}\eta(1-\bar{s}+z)}{z}dz=0$ (61)
$\Rightarrow\lim_{N\to\infty}\int_{-\infty}^{\infty}\frac{N^{i(t+y)}\eta(i(t+y))}{\sigma-1+iy}dy=0$
Applying the same procedure for $N^{s+1}\eta_{N}(s)$, we get,
$N^{s+1}\eta_{N}(s)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+1+z}\eta(s+z)}{z}dz-\frac{N}{2}$ (62)
According to equation (36)
$N^{s+1}\eta_{N}(s)=-\frac{N}{2}+N^{s+1}\eta(s)+\frac{s}{4}+O\left(\frac{1}{N}\right)=-\frac{N}{2}+\frac{s}{4}+O\left(\frac{1}{N}\right)$ (63)
Therefore,
$\lim_{N\to\infty}\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+1+z}\eta(s+z)}{z}dz=\frac{s}{4}$ (64)
Let
$A_{N}(s)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+1+z}\eta(s+z)}{z}\,dz$ (65)
By shifting the contour to $\Re\{z\}=-\sigma$ as we did before,
$\lim_{N\to\infty}A_{N}(s)=\lim_{N\to\infty}\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{s+2+iy}\eta(s+1+iy)}{1+iy}dy=\lim_{N\to\infty}\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{i(t+y)+1}\eta(i(t+y))}{-\sigma+iy}dy=\frac{s}{4}$ (66)
and similarly,
$\lim_{N\to\infty}A_{N}(1-\bar{s})=\lim_{N\to\infty}\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{3-\bar{s}+iy}\eta(2-\bar{s}+iy)}{1+iy}dy=\lim_{N\to\infty}\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{i(t+y)+1}\eta(i(t+y))}{\sigma-1+iy}dy=\frac{1-\bar{s}}{4}$ (67)
Note that,
$\lim_{N\to\infty}\mu_{N}(s)$ $=\lim_{N\to\infty}N^{s+1}\eta_{N}(s)-N^{2-\bar{s}}\eta_{N}(1-\bar{s})$ (68)
$=\lim_{N\to\infty}\left[\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+1+z}\eta(s+z)}{z}\,dz-\frac{N}{2}\right]-\left[\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{2-\bar{s}+z}\eta(1-\bar{s}+z)}{z}\,dz-\frac{N}{2}\right]$
$=\lim_{N\to\infty}[A_{N}(s)-A_{N}(1-\bar{s})]$
Let us write $\lim_{N\to\infty}\mu_{N}(s)$ as,
$\lim_{N\to\infty}\mu_{N}(s)$ $=\lim_{N\to\infty}\frac{1}{2\pi/N}\int_{-\infty}^{\infty}g_{N}(s,y)\,dy$ (69)
where
$g_{N}(s,y)=\frac{N^{i(t+y)}\eta(i(t+y))}{-\sigma+iy}-\frac{N^{i(t+y)}\eta(i(t+y)}{\sigma-1+iy}$ (70)
From equations (59) and (LABEL:eq1013), $\int_{-\infty}^{\infty}g_{N}(s,y)\,dy$ converges to 0 as $N\to\infty$. Let us investigate the convergence of $\int_{-\infty}^{\infty}\frac{\partial}{\partial N}[g_{N}(s,y)]\,dy$.
$\int_{-\infty}^{\infty}\frac{\partial}{\partial N}[g_{N}(s,y)]\,dy$ $=\int_{-\infty}^{\infty}\frac{i(t+y)N^{i(t+y)-1}\eta(i(t+y))}{-\sigma+iy}-\frac{i(t+y)N^{i(t+y)-1}\eta(i(t+y)}{\sigma-1+iy}dy$ (71)
$=\int_{-\infty}^{\infty}\left\{\frac{[(-\sigma+iy)+(\sigma+it)]N^{i(t+y)-1}\eta(i(t+y))}{-\sigma+iy}\right$.
$\quad\left.-\frac{[(\sigma-1+iy)+(1-\sigma+it)]N^{i(t+y)-1}\eta(i(t+y)}{\sigma-1+iy}\right\}dy$
$=\frac{1}{N}\int_{-\infty}^{\infty}\frac{sN^{i(t+y)}\eta(i(t+y))}{-\sigma+iy}-\frac{(1-\bar{s})N^{i(t+y)}\eta(i(t+y)}{\sigma-1+iy}dy$
This integral converges uniformly to 0 as $N\to\infty$ from equations(59) and (LABEL:eq1013).
So, by applying L’Hôpital rule and Leibniz integral rule for differentiation under the integral sign [11] to equation (69), we get
$\lim_{N\to\infty}\mu_{N}(s)$ $=\lim_{N\to\infty}[A_{N}(s)-A_{N}(1-\bar{s})]$ (72)
$=-\lim_{N\to\infty}\frac{N^{2}}{2\pi}\int_{-\infty}^{\infty}\frac{[sN^{i(t+y)-1}\eta(i(t+y))}{-\sigma+iy}-\frac{[(1-\bar{s})]N^{i(t+y)-1}\eta(i(t+y)}{\sigma-1+iy}dy$
$=-\lim_{N\to\infty}\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{[sN^{i(t+y)+1}\eta(i(t+y))}{-\sigma+iy}-\frac{[(1-\bar{s})]N^{i(t+y)+1}\eta(i(t+y)}{\sigma-1+iy}dy$
$=-\lim_{N\to\infty}[sA_{N}(s)-(1-\bar{s})A_{N}(1-\bar{s})]$
Therefore,
$\lim_{N\to\infty}(1+s)A_{N}(s)-(2-\bar{s})A_{N}(1-\bar{s})=0$ (73)
Substituting by equations (66) and (67) in equation (73),
$\lim_{N\to\infty}(1+s)A_{N}(s)-(2-\bar{s})A_{N}(1-\bar{s})$ $=\left(1+s\right)\left(\frac{s}{4}\right)-\left(2-\bar{s}\right)\left(\frac{1-\bar{s}}{4}\right)=0$ (74)
$\Rightarrow(1+\sigma+it)(\sigma+it)=(2-\sigma+it)(1-\sigma+it)$
$\Rightarrow\sigma(1+\sigma)+(1+2\sigma)it-t^{2}=(1-\sigma)(2-\sigma)+(3-2\sigma)it-t^{2}$
$\Rightarrow\sigma=1/2$
Therefore, all the non-trivial zeros of the zeta function must lie on the critical line, $\Re(s)=1/2$. ∎