### v37

# A Simple Proof of the Riemann Hypothesis

Abstract
In this article, it is proved that the non-trivial zeros of the Riemann zeta function must lie on the critical line, known as the Riemann hypothesis.
Keywords Riemann zeta function $\cdot$ Riemann hypothesis $\cdot$ Non-trivial zeros $\cdot$ Critical line
1 Riemann Zeta function
The Riemann zeta function is defined over the complex plane as [1],
$\zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n^{s}}\quad,\Re(s)>1$ (1)
where $\Re(s)$ denotes the real part of $s$. There are several forms can be used for an analytic continuation for $\Re(s)>0$ such as [1, 2],
$\zeta(s)=\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{N^{1-s}}{1-s}-s\int_{N}^{\infty}\frac{x-\lfloor x\rfloor}{x^{s+1}}dx,N\in\mathbb{N}$ (2)
where $\lfloor x\rfloor$ is the floor or integer part such that $x-1<\lfloor x\rfloor\leq x$ for real $x$.
and
$\zeta(s)=\frac{1}{1-2^{1-s}}\eta(s),\qquad\eta(s)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n^{s}}$ (3)
where $s\neq 1+\frac{2\pi ki}{\log(2)},k=0,\pm 1,\pm 2\ldots$ and $\eta(s)$ is the Dirichlet eta function (sometimes called the alternating zeta function).
Using Euler-Maclaurin formula, it can also be written as [1, 2],
$\zeta(s)=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\sum_{i=1}^{m}\binom{s+2i-2}{2i-1}\frac{B_{2i}}{2i}N^{1-s-2i}-\epsilon_{m},\quad\Re(s)>-2m;m,N\in\mathbb{N}$ (4)
where
$\zeta_{N}(s)=\sum_{n=1}^{N}\frac{1}{n^{s}},\quad\epsilon_{m}$ $=\binom{s+2m}{2m+1}\int_{N}^{\infty}\frac{\bar{B}_{2m+1}(x)}{x^{s+2m+1}}dx,$ (5)
$B_{k}$ is the $k$-th Bernoulli number defined implicitly by,
$\frac{t}{e^{t}-1}=\sum_{k=0}^{\infty}B_{k}\frac{t^{k}}{k!},$ (6)
$B_{k}(x)$ is the $k$-th Bernoulli polynomial defined as the unique polynomial of degree $k$ with the property that,
$\int_{t}^{t+1}B_{k}(x)dx=t^{k},$ (7)
$\bar{B}_{k}(x)$ is the periodic function $B_{k}(x-\lfloor x\rfloor)$.
Riemann also extended $\zeta(s)$ to $\mathbb{C}$ as a meromorphic function, with only a simple pole at $s=1$ with residue 11, by the functional equation,
$\zeta({s})=2^{s}\pi^{s-1}\sin\left(\frac{\pi s}{2}\right)\Gamma\left({1-s}\right)\zeta({1-s}),\qquad s\in\mathbb{C}\setminus\{0,1\}$ (8)
2 Zeros of the Riemann Zeta Function
The trivial zeros of the Riemann zeta function occur at the negative even integers; that is, $\zeta(-2n)=0,n\in\mathbb{N}$ [1]. On the other hand, the non-trivial zeros lie in the critical strip, $0\leq\Re(s)\leq 1$. Both Hadamard [3] and de la Vallee Poussin [4] independently proved that there are no zeros on the boundaries of the critical strip (i.e. $\Re(s)=0$ or $\Re(s)=1$). Gourdon and Demichel [5] verified the Riemann Hypothesis until the $10^{13}$-th zero.
Mossinghoff and Trudgian [6] proved that there are no zeros for $\zeta(\sigma+it)$ for $|t|\geq 2$ in the region,
$\sigma\geq 1-\frac{1}{5.573412\log|t|}$ (9)
This represents the largest known zero-free region for the zeta-function within the critical strip for $3.06\times 10^{10}<|t|<\exp(10151.5)\approx 5.5\times 10^{4408}$.
Due to the functional equation (8) and the complex conjugation properties, the non-trivial zeros of $\zeta(s)$ are symmetric with respect to both the critical line and the real axis, that is $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$. According to equation (3), the zeros of the Dirichlet eta function include all the non-trivial zeros of the zeta function. So, if $s$ is a non-trivial zero of the zeta function, then $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$.
3 Riemann Hypothesis
All the non-trivial zeros of the Riemann zeta function lie on the critical line $\Re(s)=1/2$.
Proof.
Assume that $\zeta(s)=0$ where $s=\sigma+it,0<\sigma<1,t\in\mathbb{R}$ .
Using $m=1$ in equations (4) and (5), we get,
$\zeta(s)$ $=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}-\epsilon_{1}$ (10)
and the error term is given by,
$\epsilon_{1}$ $=\binom{s+2}{3}\int_{N}^{\infty}\frac{\bar{B}_{3}(x)}{x^{s+3}}dx,$ (11)
Since,
[…]
then, the error term, $\epsilon_{1}$, can be bounded as,
$|\epsilon_{1}|$ $\leq B^{max}_{3}\frac{|s(s+1)(s+2)|}{3!}\int_{N}^{\infty}\frac{1}{x^{\sigma+3}}dx\leq B^{max}_{3}\frac{|s(s+1)(s+2)|}{(3!)(\sigma+2)}\left(\frac{1}{N^{\sigma+2}}\right)$ (13)
Therefore,
$\zeta(s)=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}+O\left(\frac{1}{N^{{}^{\sigma+2}}}\right)$ (14)
Using Perron’s formula [8],
$\zeta_{N}(s)$ $=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{N^{z}}{z}\sum_{n=1}^{\infty}\frac{1}{n^{s+z}}dz+\frac{1}{2N^{s}}=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{N^{z}\zeta(s+z)}{z}dz+\frac{1}{2N^{s}},\qquad c>1-\sigma$ (15)
Using $c=1$, multiplying by $N^{s}$,
$N^{s}\zeta_{N}(s)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\zeta(s+z)}{z}dz+\frac{1}{2}$ (16)
According to equations (14) and (16), we have,
$N^{s}\zeta_{N}(s)=N^{s}\zeta(s)+\frac{N}{1-s}+\frac{1}{2}-\frac{s}{12N}+O\left(\frac{1}{N^{2}}\right)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\zeta(s+z)}{z}dz+\frac{1}{2}$ (17)
Therefore, for $\zeta(s)=0$,
$\lim_{N\to\infty}\left\{\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\zeta(s+z)}{z}dz-\frac{N}{1-s}\right\}=\lim_{N\to\infty}\left\{\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{s+1+iy}\zeta(s+1+iy)}{1+iy}dy-\frac{N}{1-s}\right\}=0$ (18)
Let
$f_{N}(s,z)=\frac{N^{s+z}\zeta(s+z)}{z}$ (19)
Let us take a rectangular contour $C$ as,
  1. i)
    The vertical line from 1iT1-iT to $1+iT$,
  2. ii)
    The horizontal line from $1+iT$ to $1-\sigma-\delta+iT$,
  3. iii)
    The vertical line from $1-\sigma-\delta+iT$ to $1-\sigma-\delta-iT$,
  4. iv)
    The horizontal line from $1-\sigma-\delta-iT$ to 1iT1-iT.
where $T\to\infty,0<\delta<\min(\sigma,1-\sigma)$.
Note that $f_{N}(s,z)$ has a simple pole at z=1sz=1-s inside the contour $C$.
By using analytic continuation of the zeta function and applying Cauchy residue theorem [9], we get,
$\ointctrclockwise_{C}f_{N}(s,z)\,dz$ $=\int_{1-iT}^{1+iT}f_{N}(s,z)\,dz+\int_{1+iT}^{1-\sigma-\delta+iT}f_{N}(s,z)\,dz+\int_{1-\sigma-\delta+iT}^{1-\sigma-\delta-iT}f_{N}(s,z)\,dz+\int_{1-\sigma-\delta-iT}^{1-iT}f_{N}(s,z)\,dz$ (20)
$=2\pi i\operatorname{Res}_{z=1-s}f_{N}(s,z)=2\pi i\left(\frac{N}{1-s}\right)$
The second integral can be written as,
$\int_{1+iT}^{1-\sigma-\delta+iT}f_{N}(s,z)\,dz=\int_{1}^{1-\sigma-\delta}\frac{N^{s+x+iT}\zeta(s+x+iT)}{x+iT}\,dx$ (21)
and we have,
$|x+iT|\geq T$ (22)
From the convexity property of the zeta function [10], for large $T$, we have
[…]
Therefore,
$\int_{1+iT}^{1-\sigma-\delta+iT}f_{N}(s,z)\,dz\to 0\text{ as }T\to\infty$ (24)
and similarly, the fourth integral,
$\int_{1-\sigma-\delta-iT}^{1-iT}f_{N}(s,z)\,dz\to 0\text{ as }T\to\infty$ (25)
Substituting in equation (20) and taking the limit as $N\to\infty$ leads to,
$\lim_{N\to\infty}\int_{1-\sigma-\delta-i\infty}^{1-\sigma-\delta+i\infty}f_{N}(s,z)\,dz=\lim_{N\to\infty}\int_{-\infty}^{\infty}\frac{N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{1-\sigma-\delta+iy}dy=0$ (26)
Due to the symmetry of the non-trivial zeros of the zeta function about the critical line, $\zeta(1-\bar{s})=0$ and similarly, we must have,
$\lim_{N\to\infty}\int_{\sigma-\delta-i\infty}^{\sigma-\delta+i\infty}f_{N}(1-\bar{s},z)\,dz=\lim_{N\to\infty}\int_{-\infty}^{\infty}\frac{N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{\sigma-\delta+iy}dy=0$ (27)
According to equation (17), we also have,
$N^{s+1}\zeta_{N}(s)=N^{s+1}\zeta(s)+\frac{N^{2}}{1-s}+\frac{N}{2}-\frac{s}{12}+O\left(\frac{1}{N}\right)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+1+z}\zeta(s+z)}{z}dz+\frac{N}{2}$ (28)
Therefore,
$\lim_{N\to\infty}\left\{\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+1+z}\zeta(s+z)}{z}dz-N^{s+1}\zeta(s)-\frac{N^{2}}{1-s}\right\}=-\frac{s}{12}$ (29)
By shifting the contour to $\Re\{z\}=1-\sigma-\delta$ as we did before,
$\lim_{N\to\infty}\left[A_{N}(s)-N^{s+1}\zeta(s)\right]=-\frac{s}{12}$ (30)
where
$A_{N}(s)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{2-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{1-\sigma-\delta+iy}dy$ (31)
Note that $A_{N}(s)$ diverges as $N\to\infty$ for $\zeta(s)\neq 0$ due to the term $N^{s+1}\zeta(s)$ in equation (30). However, for $\zeta(s)=0$,
$\lim_{N\to\infty}A_{N}(s)=-\frac{s}{12}$ (32)
Repeating similar steps for $N^{2-\bar{s}}\zeta_{N}(1-\bar{s})$, we get
$\lim_{N\to\infty}\left[A_{N}(1-\bar{s})-N^{2-\bar{s}}\zeta(1-\bar{s})\right]=-\frac{1-\bar{s}}{12}$ (33)
and at $\zeta(1-\bar{s})=0$,
$\lim_{N\to\infty}A_{N}(1-\bar{s})=-\frac{1-\bar{s}}{12}$ (34)
and
$\lim_{N\to\infty}\left[A_{N}(s)-A_{N}(1-\bar{s})\right]=-\frac{s}{12}+\frac{1-\bar{s}}{12}=\frac{1-2\sigma}{12}$ (35)
This limit exists only for $\zeta(s)=\zeta(1-\bar{s})=0$. Let us write this limit in a different form as,
$\lim_{N\to\infty}\left[A_{N}(s)-A_{N}(1-\bar{s})\right]$ $=\lim_{N\to\infty}\frac{1}{2\pi/N}\int_{-\infty}^{\infty}g_{N}(s,y)\,dy$ (36)
where
$g_{N}(s,y)=N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]$ (37)
From equations (26) and (27), $\int_{-\infty}^{\infty}g_{N}(s,y)\,dy$ converges to 0 as $N\to\infty$. So, by extending to real values of NN and applying L’Hôpital rule [11] to equation (36), we get
$\lim_{N\to\infty}\left[A_{N}(s)-A_{N}(1-\bar{s})\right]$ $=-\lim_{N\to\infty}\frac{N^{2}}{2\pi}\lim_{h\to 0}\frac{1}{h}\int_{-\infty}^{\infty}g_{N+h}(s,y)-g_{N}(s,y)\,dy$ (38)
$=-\lim_{N\to\infty}\frac{N^{2}}{2\pi}\lim_{h\to 0}\frac{1}{h}\int_{-\infty}^{\infty}\left[(N+h)^{1-\delta+i(t+y)}-N^{1-\delta+i(t+y)}\right]\zeta(1-\delta+i(t+y))$
$\quad\times\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]\,dy$
$=-\lim_{N\to\infty}\frac{N^{2}}{2\pi}\lim_{h\to 0}\frac{1}{h}\int_{-\infty}^{\infty}N^{1-\delta+i(t+y)}\left[\left(1+\frac{h}{N}\right)^{1-\delta+i(t+y)}-1\right]\zeta(1-\delta+i(t+y))$
$\quad\times\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]\,dy$
Using the Binomial expansion for $(1+h/N)^{1-\delta+i(t+y)}$,
$\lim_{N\to\infty}\left[A_{N}(s)-A_{N}(1-\bar{s})\right]$ $=-\lim_{N\to\infty}\frac{N^{2}}{2\pi}\lim_{h\to 0}\frac{1}{h}\int_{-\infty}^{\infty}N^{1-\delta+i(t+y)}\left[\frac{(1-\delta+i(t+y))h}{N}+O\left(\frac{h^{2}}{N^{2}}\right)\right]$ (39)
$\quad\times\zeta(1-\delta+i(t+y))\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]\,dy$
$=-\frac{1}{2\pi}\lim_{N\to\infty}\lim_{h\to 0}\int_{-\infty}^{\infty}N^{2-\delta+i(t+y)}\left[(1-\delta+i(t+y))+O\left(\frac{h}{N}\right)\right]$
$\quad\times\zeta(1-\delta+i(t+y))\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]\,dy=I_{1}+I_{2}$
where
$I_{1}$ $=-\frac{1}{2\pi}\lim_{N\to\infty}\int_{-\infty}^{\infty}\left\{\frac{1-\delta+i(t+y)}{1-\sigma-\delta+iy}-\frac{1-\delta+i(t+y)}{\sigma-\delta+iy}\right\}N^{2-\delta+i(t+y)}\zeta(1-\delta+i(t+y))\,dy$ (40)
$=-\frac{1}{2\pi}\lim_{N\to\infty}\int_{-\infty}^{\infty}\left\{\frac{(1-\sigma-\delta+iy)+(\sigma+it)}{1-\sigma-\delta+iy}-\frac{(\sigma-\delta+iy)+(1-\sigma+it)}{\sigma-\delta+iy}\right\}$
$\qquad\times N^{2-\delta+i(t+y)}\zeta(1-\delta+i(t+y))\,dy$
$=-\frac{1}{2\pi}\lim_{N\to\infty}\int_{-\infty}^{\infty}\left\{\frac{s}{1-\sigma-\delta+iy}-\frac{1-\bar{s}}{\sigma-\delta+iy}\right\}N^{2-\delta+i(t+y)}\zeta(1-\delta+i(t+y))\,dy$
$=-\lim_{N\to\infty}\left[sA_{N}(s)-(1-\bar{s})A_{N}(1-\bar{s})\right]$
and
$I_{2}=-\frac{1}{2\pi}\lim_{N\to\infty}\lim_{h\to 0}\int_{-\infty}^{\infty}N^{2-\delta+i(t+y)}O\left(\frac{h}{N}\right)\zeta(1-\delta+i(t+y))\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]\,dy$ (41)
Since $I_{1}$ converges according to equations (32) and (34), then $I_{2}$ must converge too. Thus, the limit in $I_{2}$ must exist, and accordingly the integral in $I_{2}$ must converge. This yields $I_{2}=0$ because of the $O(h)$ factor .
Therefore,
$\lim_{N\to\infty}\left[A_{N}(s)-A_{N}(1-\bar{s})\right]$ $=-\lim_{N\to\infty}\left[sA_{N}(s)-(1-\bar{s})A_{N}(1-\bar{s})\right]$ (42)
That is,
$\lim_{N\to\infty}(1+s)A_{N}(s)-(2-\bar{s})A_{N}(1-\bar{s})=0$ (43)
Substituting by equations (32) and (34) in equation (43),
$\lim_{N\to\infty}(1+s)A_{N}(s)-(2-\bar{s})A_{N}(1-\bar{s})$ $=\left(1+s\right)\left(-\frac{s}{12}\right)-\left(2-\bar{s}\right)\left(-\frac{1-\bar{s}}{12}\right)=0$ (44)
$s(1+s)$ $=(1-\bar{s})(2-\bar{s})$
$\Rightarrow(\sigma+it)(1+\sigma+it)$ $=(1-\sigma+it)(2-\sigma+it)$
$\Rightarrow\sigma(1+\sigma)+(1+2\sigma)it-t^{2}$ $=(1-\sigma)(2-\sigma)+(3-2\sigma)it-t^{2}$
$\Rightarrow\sigma$ $=1/2$
Therefore, all the non-trivial zeros of the zeta function must lie on the critical line, $\Re(s)=1/2$. ∎