### v39

# A Simple Proof of the Riemann Hypothesis

Abstract
In this article, it is proved that the non-trivial zeros of the Riemann zeta function must lie on the critical line, known as the Riemann hypothesis.
Keywords Riemann zeta function $\cdot$ Riemann hypothesis $\cdot$ Non-trivial zeros $\cdot$ Critical line
1 Riemann Zeta function
The Riemann zeta function is defined over the complex plane as [1],
$\zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n^{s}}\quad,\Re(s)>1$ (1)
where $\Re(s)$ denotes the real part of $s$. There are several forms can be used for an analytic continuation for $\Re(s)>0$ such as [1, 2],
$\zeta(s)=\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{N^{1-s}}{1-s}-s\int_{N}^{\infty}\frac{x-\lfloor x\rfloor}{x^{s+1}}dx,N\in\mathbb{N}$ (2)
where $\lfloor x\rfloor$ is the floor or integer part such that $x-1<\lfloor x\rfloor\leq x$ for real $x$.
and
$\zeta(s)=\frac{1}{1-2^{1-s}}\eta(s),\qquad\eta(s)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n^{s}}$ (3)
where $s\neq 1+\frac{2\pi ki}{\log(2)},k=0,\pm 1,\pm 2\ldots$ and $\eta(s)$ is the Dirichlet eta function (sometimes called the alternating zeta function).
Using Euler-Maclaurin formula, it can also be written as [1, 2],
$\zeta(s)=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\sum_{i=1}^{m}\binom{s+2i-2}{2i-1}\frac{B_{2i}}{2i}N^{1-s-2i}-\epsilon_{m},\quad\Re(s)>-2m;m,N\in\mathbb{N}$ (4)
where
$\zeta_{N}(s)=\sum_{n=1}^{N}\frac{1}{n^{s}},\quad\epsilon_{m}$ $=\binom{s+2m}{2m+1}\int_{N}^{\infty}\frac{\bar{B}_{2m+1}(x)}{x^{s+2m+1}}dx,$ (5)
$B_{k}$ is the $k$-th Bernoulli number defined implicitly by,
$\frac{t}{e^{t}-1}=\sum_{k=0}^{\infty}B_{k}\frac{t^{k}}{k!},$ (6)
$B_{k}(x)$ is the $k$-th Bernoulli polynomial defined as the unique polynomial of degree $k$ with the property that,
$\int_{t}^{t+1}B_{k}(x)dx=t^{k},$ (7)
$\bar{B}_{k}(x)$ is the periodic function $B_{k}(x-\lfloor x\rfloor)$.
Riemann also extended $\zeta(s)$ to $\mathbb{C}$ as a meromorphic function, with only a simple pole at $s=1$ with residue 11, by the functional equation,
$\zeta({s})=2^{s}\pi^{s-1}\sin\left(\frac{\pi s}{2}\right)\Gamma\left({1-s}\right)\zeta({1-s}),\qquad s\in\mathbb{C}\setminus\{0,1\}$ (8)
2 Zeros of the Riemann Zeta Function
The trivial zeros of the Riemann zeta function occur at the negative even integers; that is, $\zeta(-2n)=0,n\in\mathbb{N}$ [1]. On the other hand, the non-trivial zeros lie in the critical strip, $0\leq\Re(s)\leq 1$. Both Hadamard [3] and de la Vallee Poussin [4] independently proved that there are no zeros on the boundaries of the critical strip (i.e. $\Re(s)=0$ or $\Re(s)=1$). Gourdon and Demichel [5] verified the Riemann Hypothesis until the $10^{13}$-th zero.
Mossinghoff and Trudgian [6] proved that there are no zeros for $\zeta(\sigma+it)$ for $|t|\geq 2$ in the region,
$\sigma\geq 1-\frac{1}{5.573412\log|t|}$ (9)
This represents the largest known zero-free region for the zeta-function within the critical strip for $3.06\times 10^{10}<|t|<\exp(10151.5)\approx 5.5\times 10^{4408}$.
Due to the functional equation (8) and the complex conjugation properties, the non-trivial zeros of $\zeta(s)$ are symmetric with respect to both the critical line and the real axis, that is $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$. According to equation (3), the zeros of the Dirichlet eta function include all the non-trivial zeros of the zeta function. So, if $s$ is a non-trivial zero of the zeta function, then $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$.
3 Riemann Hypothesis
All the non-trivial zeros of the Riemann zeta function lie on the critical line $\Re(s)=1/2$.
Proof.
Assume that $s=\sigma+it,0<\sigma<1,t\in\mathbb{R}$ . Using $m=1$ in equations (4) and (5), we get,
$\zeta(s)$ $=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}-\epsilon_{1}$ (10)
and the error term is given by,
$\epsilon_{1}$ $=\binom{s+2}{3}\int_{N}^{\infty}\frac{\bar{B}_{3}(x)}{x^{s+3}}dx,$ (11)
Since,
[…]
then, the error term, $\epsilon_{1}$, can be bounded as,
$|\epsilon_{1}|$ $\leq B^{max}_{3}\frac{|s(s+1)(s+2)|}{3!}\int_{N}^{\infty}\frac{1}{x^{\sigma+3}}dx\leq B^{max}_{3}\frac{|s(s+1)(s+2)|}{(3!)(\sigma+2)}\left(\frac{1}{N^{\sigma+2}}\right)$ (13)
Therefore,
$\zeta(s)=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\frac{s}{12N^{s+1}}+O\left(\frac{1}{N^{{}^{\sigma+2}}}\right)$ (14)
Multiplying by $N^{s}$, we get
$N^{s}\zeta(s)=N^{s}\zeta_{N}(s)-\frac{N}{1-s}-\frac{1}{2}+O\left(\frac{1}{N}\right)$ (15)
Using Perron’s formula [8],
$\zeta_{N}(s)$ $=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{N^{z}}{z}\sum_{n=1}^{\infty}\frac{1}{n^{s+z}}dz+\frac{1}{2N^{s}}=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{N^{z}\zeta(s+z)}{z}dz+\frac{1}{2N^{s}},\qquad c>1-\sigma$ (16)
Using $c=1$ and multiplying by $N^{s}$,
$N^{s}\zeta_{N}(s)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\zeta(s+z)}{z}dz+\frac{1}{2}$ (17)
According to equations (15) and (17), we have,
$N^{s}\zeta_{N}(s)=N^{s}\zeta(s)+\frac{N}{1-s}+\frac{1}{2}+O\left(\frac{1}{N}\right)=\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\zeta(s+z)}{z}dz+\frac{1}{2}$ (18)
Therefore,
$\frac{1}{2\pi i}\int_{1-i\infty}^{1+i\infty}\frac{N^{s+z}\zeta(s+z)}{z}dz$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{1+\sigma+i(t+y)}\zeta(1+\sigma+i(t+y))}{1+iy}dy=N^{s}\zeta(s)+\frac{N}{1-s}+O\left(\frac{1}{N}\right)$ (19)
Let
$f_{N}(s,z)=\frac{N^{s+z}\zeta(s+z)}{z}$ (20)
Let us take a rectangular contour $C$ as,
  1. i)
    The vertical line from 1iT1-iT to $1+iT$,
  2. ii)
    The horizontal line from $1+iT$ to $1-\sigma-\delta+iT$,
  3. iii)
    The vertical line from $1-\sigma-\delta+iT$ to $1-\sigma-\delta-iT$,
  4. iv)
    The horizontal line from $1-\sigma-\delta-iT$ to 1iT1-iT.
where $T\to\infty,0<\delta<\min(\sigma,1-\sigma)$ (see Figure 1).
Figure 1: Contour $C$ used to shift the line of integration from $\Re\{z\}=1$ to $\Re\{z\}=1-\sigma-\delta$.
Note that $f_{N}(s,z)$ has a simple pole at z=1sz=1-s inside the contour $C$.
By using analytic continuation of the zeta function and applying Cauchy residue theorem [9], we get,
$\ointctrclockwise_{C}f_{N}(s,z)\,dz$ $=\int_{1-iT}^{1+iT}f_{N}(s,z)\,dz+\int_{1+iT}^{1-\sigma-\delta+iT}f_{N}(s,z)\,dz+\int_{1-\sigma-\delta+iT}^{1-\sigma-\delta-iT}f_{N}(s,z)\,dz+\int_{1-\sigma-\delta-iT}^{1-iT}f_{N}(s,z)\,dz$ (21)
$=2\pi i\operatorname{Res}_{z=1-s}f_{N}(s,z)=2\pi i\left(\frac{N}{1-s}\right)$
The second integral can be written as,
$\int_{1+iT}^{1-\sigma-\delta+iT}f_{N}(s,z)\,dz=\int_{1}^{1-\sigma-\delta}\frac{N^{s+x+iT}\zeta(s+x+iT)}{x+iT}\,dx$ (22)
We have,
$|x+iT|\geq T$ (23)
and from the convexity property of the zeta function [10], for large $T$,
[…]
where $\varepsilon>0$ is an arbitrarily small positive constant, and the implied constant in the $\ll_{\varepsilon}$-notation depends on $\varepsilon$.
Therefore, for a fixed NN,
$\int_{1+iT}^{1-\sigma-\delta+iT}f_{N}(s,z)\,dz\to 0\text{ as }T\to\infty$ (25)
and similarly, the fourth integral,
$\int_{1-\sigma-\delta-iT}^{1-iT}f_{N}(s,z)\,dz\to 0\text{ as }T\to\infty$ (26)
Substituting by equations (21), (25) and (26) in equation (19), we get
$\frac{1}{2\pi i}\int_{1-\sigma-\delta-i\infty}^{1-\sigma-\delta+i\infty}f_{N}(s,z)\,dz=\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{1-\sigma-\delta+iy}dy=N^{s}\zeta(s)+O\left(\frac{1}{N}\right)$ (27)
Similarly, for $\zeta(1-\bar{s})=0$, by replacing $s$ by $1-\bar{s}$ (i.e. $\sigma$ by $1-\sigma$), we get
$\frac{1}{2\pi i}\int_{\sigma-\delta-i\infty}^{\sigma-\delta+i\infty}f_{N}(1-\bar{s},z)\,dz=\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{\sigma-\delta+iy}dy=N^{1-\bar{s}}\zeta(1-\bar{s})+O\left(\frac{1}{N}\right)$ (28)
Subtracting equation (28) from equation (27), we get
$\frac{1}{2\pi}\int_{-\infty}^{\infty}\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))dy=N^{s}\zeta(s)-N^{1-\bar{s}}\zeta(1-\bar{s})+O\left(\frac{1}{N}\right)$ (29)
or
$\frac{(2\sigma-1)}{2\pi}\int_{-\infty}^{\infty}\frac{N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{\left(1-\sigma-\delta+iy\right)\left(\sigma-\delta+iy\right)}dy=N^{s}\zeta(s)-N^{1-\bar{s}}\zeta(1-\bar{s})+O\left(\frac{1}{N}\right)$ (30)
Let us study the following two cases for this equation.
Case I: $\sigma=1/2$
In this case, both sides are identically zero. So, the equation vanishes at all points along the critical line, $\sigma=1/2$, including the well-known non-trivial zeros, $\zeta(s)=0$, and the non-zeros, $\zeta(s)\neq 0$.
Case II: $\sigma\neq 1/2$
Equation (LABEL:eq120) can be written as,
$\int_{-\infty}^{\infty}\frac{N^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{\left(1-\sigma-\delta+iy\right)\left(\sigma-\delta+iy\right)}dy=\frac{2\pi}{2\sigma-1}\left[N^{s}\zeta(s)-N^{1-\bar{s}}\zeta(1-\bar{s})\right]+O\left(\frac{1}{N}\right)$ (31)
As $N\to\infty$, the right-hand side either
  • diverges if $\zeta(s)\neq 0$, and accordingly, $\zeta(1-\bar{s})\neq 0$, as either $N^{s}\zeta(s)$ or $N^{1-\bar{s}}\zeta(1-\bar{s})$ will dominate.
  • converges to 0 if $\zeta(s)=\zeta(1-\bar{s})=0$.
In the sequel, we will show that the left-hand side is always divergent and can never converge to 0 for $\sigma\neq 1/2$ ensuring that $\zeta(s)$ cannot be zero in this region.
The left-hand side of equation (LABEL:eq121) can be written as,
$I_{N}(s)=N^{1-\delta+it}\int_{-\infty}^{\infty}g(y)e^{iy\log N}dy$ (32)
where
$g(y)=\frac{\zeta(1-\delta+i(t+y))}{\left(1-\sigma-\delta+iy\right)\left(\sigma-\delta+iy\right)}$ (33)
It can also be viewed as,
$I_{N}(s)=N^{1-\delta+it}\hat{g}(-\log N)$ (34)
where $\hat{g}(\xi)$ is the Fourier transform of $g(y)$ given by,
$\hat{g}(\xi)=\int_{-\infty}^{\infty}{g}(y)e^{-iy\xi}dy$ (35)
For $A>0$, an arbitrary positive constant (independent of NN), $r(N)=A/\log N$ and $w(N)=A/\sqrt{\log N}$, let $\phi_{N}(y)$ be a smooth cutoff function satisfying $\phi_{N}(y)=1$ on the near-field region $|y|\leq r(N)$, where the oscillatory phase $y\log N$ varies slowly, and $\phi_{N}(y)=0$ for $|y|\geq r(N)+w(N)$. The transition occurs smoothly in the annulus $r(N)\leq|y|\leq r(N)+w(N)$. We then decompose the integral $I_{N}(s)$ as,
$I_{N}(s)$ $=N^{1-\delta+it}\left[\int_{|y|\leq r(N)}g(y)e^{iy\log N}dy+\int_{\mathbb{R}}\left[1-\phi_{N}(y)\right]g(y)e^{iy\log N}dy\right]=N^{1-\delta+it}[J_{N}(s)+H_{N}(s)]$ (36)
By substituting by $u=y\log N$ in $J_{N}(s)$, we get,
$J_{N}(s)=\frac{1}{\log N}\int_{|u|\leq A}g\left(\frac{u}{\log N}\right)e^{iu}du$ (37)
Using Taylor series,
$g\left(\frac{u}{\log N}\right)=g(0)+O\left(\frac{u}{\log N}\right)$ (38)
where
$g(0)=\frac{\zeta(1-\delta+it)}{(1-\sigma-\delta)(\sigma-\delta)}$ (39)
Thus,
$J_{N}(s)=\frac{g(0)}{\log N}\int_{-A}^{A}e^{iu}du+O\left(\frac{1}{(\log N)^{2}}\right)=\frac{2g(0)\sin A}{\log N}+O\left(\frac{1}{(\log N)^{2}}\right)$ (40)
On the other hand, let us write $H_{N}(s)$ as,
$H_{N}(s)=\int_{\mathbb{R}}g_{N}(y)e^{iy\log N}dy$ (41)
where $g_{N}(y)=\left[1-\phi_{N}(y)\right]g(y)$.
Integrating by parts twice yields,
$H_{N}(s)=\frac{1}{(i\log N)^{2}}\int_{\mathbb{R}}g^{\prime\prime}_{N}(y)e^{iy\log N}dy$ (42)
Hence,
$\left|H_{N}(s)\right|\leq\frac{\|g_{N}^{\prime\prime}(y)\|_{1}}{(\log N)^{2}}$ (43)
where $\|g_{N}^{\prime\prime}(y)\|_{1}$ is the $L^{1}$-norm of the second derivative given by,
$\|g_{N}^{\prime\prime}(y)\|_{1}$ $=\int_{\mathbb{R}}|g_{N}^{\prime\prime}(y)|dy=\int_{\mathbb{R}}\left|\left[1-\phi_{N}(y)\right]g^{\prime\prime}(y)-2\phi^{\prime}_{N}(y)g^{\prime}(y)-\phi^{\prime\prime}_{N}(y)g(y)\right|dy$ (44)
which can be bounded by
$\|g_{N}^{\prime\prime}(y)\|_{1}\leq\|\left[1-\phi_{N}(y)\right]g^{\prime\prime}(y)\|_{1}+2\|\phi^{\prime}_{N}(y)g^{\prime}(y)\|_{1}+\|\phi^{\prime\prime}_{N}(y)g(y)\|_{1}$ (45)
For large $|y|$, we have,
$\left|\frac{1}{(1-\sigma-\delta+iy)(\sigma-\delta+iy)}\right|=O\left(\frac{1}{y^{2}}\right)$ (46)
and from the convexity property of the zeta function [10],
$\left|\zeta(1-\delta+i(t+y)\right|$ $\ll_{\varepsilon}O\left(|y|^{\delta/2+\varepsilon}\right),$ (47)
$\left|\zeta^{(m)}(1-\delta+i(t+y)\right|$ $\ll_{\varepsilon}O\left(|y|^{\delta/2+\varepsilon}(\log|y|)^{m}\right),\qquad m\in\mathbb{N}$
This yields,
$g,g^{\prime},g^{\prime\prime}\in L^{1}(\mathbb{R})$ (48)
where $L^{1}(\mathbb{R})$ denotes the space of absolutely integrable functions on the real line.
Therefore,
$\|\left[1-\phi_{N}(y)\right]g^{\prime\prime}(y)\|_{1}\leq\int_{|y|\geq r(N)}\left|g^{\prime\prime}(y)\right|dy=O(1)$ (49)
Since $\max|\phi^{\prime}_{N}(y)|=O(1/w(N))$ and $\max|\phi^{\prime\prime}_{N}(y)|=O(1/(w(N))^{2})$, then we have,
$\|\phi^{\prime}_{N}(y)g^{\prime}(y)\|_{1}\leq\max|\phi^{\prime}_{N}(y)|\int_{r(N)\leq|y|\leq r(N)+w(N)}\left|g^{\prime}(y)\right|dy\leq O\left(\frac{1}{w(N)}\right)O\left(w(N)\right)=O(1),$ (50)
$\|\phi^{\prime\prime}_{N}(y)g(y)\|_{1}\leq\max|\phi^{\prime\prime}_{N}(y)|\int_{r(N)\leq|y|\leq r(N)+w(N)}\left|g(y)\right|dy\leq O\left(\frac{1}{(w(N))^{2}}\right)O\left(w(N)\right)=O\left(\frac{1}{w(N)}\right)$ (51)
Thus,
$\|g_{N}^{\prime\prime}(y)\|_{1}=O\left(\frac{1}{w(N)}\right)=O\left(\sqrt{\log N}\right)$ (52)
Substituting by equation (52) in equation (43) leads to,
$|H_{N}(s)|=O\left(\frac{1}{(\log N)^{3/2}}\right)$ (53)
For each $\zeta(\sigma+it)=\zeta(1-\sigma+it)=0$, we can always choose $0<\delta\ll\min(\sigma,1-\sigma)$ such that at the same $t$, $(1-\delta+it)$ lies in the zero-free region of the zeta function (guaranteed by [6], [5]), that is $\zeta(1-\delta+it)\neq 0$, and accordingly $g(0)\neq 0$, and $A\neq k\pi,k\in\mathbb{N}$ so $\sin A\neq 0$. Consequently, from equations (40) and (53), we have
$|I_{N}(s)|\gtrsim\left|\frac{2N^{1-\delta+it}g(0)\sin A}{\log N}\right|=\frac{2N^{1-\delta}\left|g(0)\sin A\right|}{\log N}$ (54)
So, as NN increases, the modulus of the left-hand side of equation (LABEL:eq121) blows up while that of the right-hand side diminishes, leading to a contradiction, if we assume that $\zeta(s)$ can vanish for $\sigma\neq 1/2$. Hence, $\zeta(s)$ cannot be zero for $\sigma\neq 1/2$, that is, all the non-trivial zeros of the zeta function must lie on the critical line, $\Re(s)=1/2$, concluding the proof of the Riemann hypothesis. ∎