### v42

# A Simple Proof of the Riemann Hypothesis

Abstract
In this article, it is proved that the non-trivial zeros of the Riemann zeta function must lie on the critical line, known as the Riemann hypothesis.
Keywords Riemann zeta function $\cdot$ Riemann hypothesis $\cdot$ Non-trivial zeros $\cdot$ Critical line
1 Riemann Zeta function
The Riemann zeta function is defined over the complex plane as [Riemann1859],
$\zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n^{s}}\quad,\Re(s)>1$ (1)
where $\Re(s)$ denotes the real part of $s$. There are several forms can be used for an analytic continuation for $\Re(s)>0$ such as [Riemann1859, olver2010],
$\zeta(s)=\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{N^{1-s}}{1-s}-s\int_{N}^{\infty}\frac{x-\lfloor x\rfloor}{x^{s+1}}dx,N\in\mathbb{N}$ (2)
where $\lfloor x\rfloor$ is the floor or integer part such that $x-1<\lfloor x\rfloor\leq x$ for real $x$.
and
$\zeta(s)=\frac{1}{1-2^{1-s}}\eta(s),\qquad\eta(s)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{n^{s}}$ (3)
where $s\neq 1+\frac{2\pi ki}{\log(2)},k=0,\pm 1,\pm 2\ldots$ and $\eta(s)$ is the Dirichlet eta function (sometimes called the alternating zeta function).
Using Euler-Maclaurin formula, it can also be written as [Riemann1859, olver2010],
$\zeta(s)=\zeta_{N}(s)-\frac{N^{1-s}}{1-s}-\frac{1}{2N^{s}}+\sum_{i=1}^{m}\binom{s+2i-2}{2i-1}\frac{B_{2i}}{2i}N^{1-s-2i}-\epsilon_{m},\quad\Re(s)>-2m;m,N\in\mathbb{N}$ (4)
where
$\zeta_{N}(s)=\sum_{n=1}^{N}\frac{1}{n^{s}},\quad\epsilon_{m}$ $=\binom{s+2m}{2m+1}\int_{N}^{\infty}\frac{\bar{B}_{2m+1}(x)}{x^{s+2m+1}}dx,$ (5)
$B_{k}$ is the $k$-th Bernoulli number defined implicitly by,
$\frac{t}{e^{t}-1}=\sum_{k=0}^{\infty}B_{k}\frac{t^{k}}{k!},$ (6)
$B_{k}(x)$ is the $k$-th Bernoulli polynomial defined as the unique polynomial of degree $k$ with the property that,
$\int_{t}^{t+1}B_{k}(x)dx=t^{k},$ (7)
$\bar{B}_{k}(x)$ is the periodic function $B_{k}(x-\lfloor x\rfloor)$.
Riemann also extended $\zeta(s)$ to $\mathbb{C}$ as a meromorphic function, with only a simple pole at $s=1$ with residue 11, by the functional equation,
$\zeta({s})=2^{s}\pi^{s-1}\sin\left(\frac{\pi s}{2}\right)\Gamma\left({1-s}\right)\zeta({1-s}),\qquad s\in\mathbb{C}\setminus\{0,1\}$ (8)
2 Zeros of the Riemann Zeta Function
The trivial zeros of the Riemann zeta function occur at the negative even integers; that is, $\zeta(-2n)=0,n\in\mathbb{N}$ [Riemann1859]. On the other hand, the non-trivial zeros lie in the critical strip, $0\leq\Re(s)\leq 1$. Both Hadamard [Hadamard1896] and de la Vallee Poussin [Vallee-Poussin1896] independently proved that there are no zeros on the boundaries of the critical strip (i.e. $\Re(s)=0$ or $\Re(s)=1$). Gourdon and Demichel [Gourdon2004] verified the Riemann Hypothesis until the $10^{13}$-th zero.
Mossinghoff and Trudgian [Mossinghoff2015] proved that there are no zeros for $\zeta(\sigma+it)$ for $|t|\geq 2$ in the region,
$\sigma\geq 1-\frac{1}{5.573412\log|t|}$ (9)
This represents the largest known zero-free region for the zeta-function within the critical strip for $3.06\times 10^{10}<|t|<\exp(10151.5)\approx 5.5\times 10^{4408}$.
Due to the functional equation (8) and the complex conjugation properties, the non-trivial zeros of $\zeta(s)$ are symmetric with respect to both the critical line and the real axis, that is $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$. According to equation (3), the zeros of the Dirichlet eta function include all the non-trivial zeros of the zeta function. So, if $s$ is a non-trivial zero of the zeta function, then $\zeta(s)=\zeta(1-s)=\zeta(\bar{s})=\zeta(1-\bar{s})=0$.
3 Riemann Hypothesis
All the non-trivial zeros of the Riemann zeta function lie on the critical line $\Re(s)=1/2$.
Proof.
Assume that $s=\sigma+it,0<\sigma<1,t\in\mathbb{R}$. For sufficiently large $u\in\mathbb{R^{+}}$, using the general Euler-Maclaurin formula, we have [ivic2003],
$\sum_{n\leq u}\frac{1}{n^{s}}=\zeta(s)+\frac{u^{1-s}}{1-s}-\frac{\psi(u)}{u^{s}}-\frac{sD(u)}{u^{s+1}}+O\left(\frac{1}{u^{2+\sigma}}\right)$ (10)
where
$\psi(u)$ $=u-\lfloor u\rfloor-\frac{1}{2},\qquad D(u)$ $=\frac{1}{2}\psi^{2}(u)-\frac{1}{24}$ (11)
Multiplying by $u^{s}$, we get
$\sum_{n\leq u}\left(\frac{u}{n}\right)^{s}=u^{s}\zeta(s)+\frac{u}{1-s}-\psi(u)-\frac{sD(u)}{u}+O\left(\frac{1}{u^{2}}\right)$ (12)
Using Perron’s formula [Apostol1976], for sufficiently large $T\in\mathbb{R^{+}}$ and $T\gg|t|$,
$\sideset{}{{}^{*}}{\sum}_{n\leq u}\frac{1}{n^{s}}$ $=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{u^{z}}{z}\sum_{n=1}^{\infty}\frac{1}{n^{s+z}}dz+R(u,s,T,c)=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{u^{z}\zeta(s+z)}{z}dz+R(u,s,T,c),\quad c>1-\sigma$ (13)
where
$|R(u,s,T,c)|\leq\frac{2}{T}\frac{u^{c}}{\log\left(\frac{\lfloor u\rfloor+1}{u}\right)}\sum_{n>u}\frac{1}{n^{\sigma+c}}$ (14)
and $\sum^{*}$ means that the last term in the sum must be multiplied by $1/2$ when $u$ is an integer.
Using $c=2$ and multiplying by $u^{s}$,
$\sideset{}{{}^{*}}{\sum}_{n\leq u}\left(\frac{u}{n}\right)^{s}$ $=\frac{1}{2\pi i}\int_{2-iT}^{2+iT}\frac{u^{s+z}\zeta(s+z)}{z}dz+O\left(\frac{u^{2+\sigma}}{T}\right)$ (15)
Let
$f_{u}(s,z)=\frac{u^{s+z}\zeta(s+z)}{z}$ (16)
Let us take a rectangular contour $C$ as,
  1. i)
    The vertical line from 2iT2-iT to $2+iT$,
  2. ii)
    The horizontal line from $2+iT$ to $1-\sigma-\delta+iT$,
  3. iii)
    The vertical line from $1-\sigma-\delta+iT$ to $1-\sigma-\delta-iT$,
  4. iv)
    The horizontal line from $1-\sigma-\delta-iT$ to 2iT2-iT.
where $0<\delta<\min(\sigma,1-\sigma)$ (see Figure 1).
Figure 1: Contour $C$ used to shift the line of integration from $\Re\{z\}=2$ to $\Re\{z\}=1-\sigma-\delta$.
Note that $f_{u}(s,z)$ has a simple pole at z=1sz=1-s inside the contour $C$.
By using analytic continuation of the zeta function and applying Cauchy residue theorem [stein2003], we get,
$\ointctrclockwise_{C}f_{u}(s,z)\,dz$ $=\int_{2-iT}^{2+iT}f_{u}(s,z)\,dz+\int_{2+iT}^{1-\sigma-\delta+iT}f_{u}(s,z)\,dz+\int_{1-\sigma-\delta+iT}^{1-\sigma-\delta-iT}f_{u}(s,z)\,dz+\int_{1-\sigma-\delta-iT}^{2-iT}f_{u}(s,z)\,dz$ (17)
$=2\pi i\operatorname{Res}_{z=1-s}f_{u}(s,z)=2\pi i\left(\frac{u}{1-s}\right)$
The second integral can be written as,
$\int_{2+iT}^{1-\sigma-\delta+iT}f_{u}(s,z)\,dz=\int_{2}^{1-\sigma-\delta}\frac{u^{s+x+iT}\zeta(s+x+iT)}{x+iT}\,dx$ (18)
We have,
$|x+iT|\geq T$ (19)
and from the convexity property of the zeta function [titchmarsh1986], for large $T$,
[…]
where $\varepsilon>0$ is an arbitrarily small positive constant, and the implied constant in $O(\cdot)$ depends on $\varepsilon$.
Therefore,
$\left|\int_{2+iT}^{1-\sigma-\delta+iT}f_{u}(s,z)\,dz\right|=O\left(\frac{u^{2+\sigma}}{T^{1-\delta/2-\varepsilon}}\right)$ (21)
and similarly for the fourth integral.
Substituting in equation (15) and setting $T=u^{6}$, we get
$\sideset{}{{}^{*}}{\sum}_{n\leq u}\left(\frac{u}{n}\right)^{s}$ $=\frac{1}{2\pi}\int_{-u^{6}}^{u^{6}}\frac{u^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{1-\sigma-\delta+iy}dy+\frac{u}{1-s}+O\left(\frac{1}{u^{4-\sigma}}\right)+O\left(\frac{1}{u^{4-\sigma-3\delta-6\varepsilon}}\right)$ (22)
Note that the left-hand side can be written as,
[…]
Thus, according to equations (12), (22) and (23), we have
$\sum_{n\leq u}\left(\frac{u}{n}\right)^{s}$ $=u^{s}\zeta(s)+\frac{u}{1-s}-\psi(u)-\frac{sD(u)}{u}+O\left(\frac{1}{u^{2}}\right)$ (24)
[…]
Therefore,
$A_{u}(s)$ $=\int_{-u^{6}}^{u^{6}}\frac{u^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{1-\sigma-\delta+iy}dy$ (25)
$=2\pi\left[u^{s}\zeta(s)-\psi(u)-\frac{sD(u)}{u}+O\left(\frac{1}{u^{2}}\right)-\begin{cases}1/2&u\in\mathbb{N}\\ 0&u\notin\mathbb{N}\\ \end{cases}\right]$
Similarly,
$A_{u}(1-\bar{s})$ $=\int_{-u^{6}}^{u^{6}}\frac{u^{1-\delta+i(t+y)}\zeta(1-\delta+i(t+y))}{\sigma-\delta+iy}dy$ (26)
$=2\pi\left[u^{1-\bar{s}}\zeta(1-\bar{s})-\psi(u)-\frac{(1-\bar{s})D(u)}{u}+O\left(\frac{1}{u^{2}}\right)-\begin{cases}1/2&u\in\mathbb{N}\\ 0&u\notin\mathbb{N}\\ \end{cases}\right]$
Subtracting and taking the limit as $u\to\infty$, we get,
$\lim_{u\to\infty}I_{u}(s)=\lim_{u\to\infty}[A_{u}(s)-A_{u}(1-\bar{s})]=2\pi\lim_{u\to\infty}\left[u^{s}\zeta(s)-u^{1-\bar{s}}\zeta(1-\bar{s})-(2\sigma-1)\frac{D(u)}{u}+O\left(\frac{1}{u^{2}}\right)\right]$ (27)
Let us study the following two cases for this equation.
Case I: $\sigma=1/2$
In this case, both sides are identically zero. So, the equation is satisfied at all points along the critical line, $\sigma=1/2$, including the well-known non-trivial zeros, $\zeta(s)=0$, and the non-zeros, $\zeta(s)\neq 0$.
Case II: $\sigma\neq 1/2$
Assume that $\zeta(s)=\zeta(1-\bar{s})=0$. Then, the right-hand side approaches zero uniformly in $u$ since the first two terms will be identically zero, $|\psi(u)|<1/2$ and hence $|(2\sigma-1)D(u)/u|<1/(12u)$ and so forth for the subsequent terms.
Therefore, the left-hand side must approach zero uniformly in $u$ too, that is,
$\lim_{u\to\infty}I_{u}(s)=\lim_{u\to\infty}u^{1-\delta+it}\int_{-u^{6}}^{u^{6}}g(y)dy=0$ (28)
where
$g(y)$ $=\left[\frac{1}{1-\sigma-\delta+iy}-\frac{1}{\sigma-\delta+iy}\right]u^{iy}\zeta(1-\delta+i(t+y))$ (29)
$=\left[\frac{2\sigma-1}{(1-\sigma-\delta+iy)(\sigma-\delta+iy)}\right]e^{iy\log u}\zeta(1-\delta+i(t+y))$
For $A>0$, an arbitrary positive constant (independent of $u$), $r(u)=A/\log u$ and $w(u)=A/\sqrt{\log u}$, let $\phi_{u}(y)$ be a smooth cutoff function satisfying $\phi_{u}(y)=1$ on the near-field region $|y|\leq r(u)$, where the oscillatory phase $y\log u$ varies slowly, and $\phi_{u}(y)=0$ for $|y|\geq r(u)+w(u)$. The transition occurs smoothly in the annulus $r(u)\leq|y|\leq r(u)+w(u)$. We then decompose the integral $I_{u}(s)$ as,
$I_{u}(s)$ $=u^{1-\delta+it}\left[\int_{|y|\leq r(u)}g(y)e^{iy\log u}dy+\int_{-u^{6}}^{u^{6}}\left[1-\phi_{u}(y)\right]g(y)e^{iy\log u}dy\right]=u^{1-\delta+it}[J_{u}(s)+H_{u}(s)]$ (30)
By substituting by $v=y\log u$ in $J_{u}(s)$, we get,
$J_{u}(s)=\frac{1}{\log u}\int_{|v|\leq A}g\left(\frac{v}{\log u}\right)e^{iv}dv$ (31)
Using Taylor series,
$g\left(\frac{v}{\log u}\right)=g(0)+O\left(\frac{v}{\log u}\right)$ (32)
where
$g(0)=\frac{(2\sigma-1)\zeta(1-\delta+it)}{(1-\sigma-\delta)(\sigma-\delta)}$ (33)
Thus,
$J_{u}(s)=\frac{g(0)}{\log u}\int_{-A}^{A}e^{iv}dv+O\left(\frac{1}{(\log u)^{2}}\right)=\frac{2g(0)\sin A}{\log u}+O\left(\frac{1}{(\log u)^{2}}\right)$ (34)
On the other hand, let us write $H_{u}(s)$ as,
$H_{u}(s)=\int_{-u^{6}}^{u^{6}}g_{u}(y)e^{iy\log u}dy$ (35)
where $g_{u}(y)=\left[1-\phi_{u}(y)\right]g(y)$.
Integrating by parts twice yields,
$H_{u}(s)$ $=\frac{1}{(i\log u)^{2}}\int_{-u^{6}}^{u^{6}}g^{\prime\prime}_{u}(y)e^{iy\log u}dy+\left[\frac{g_{u}(y)e^{iy\log u}}{i\log u}+\frac{g^{\prime}_{u}(y)e^{iy\log u}}{(\log u)^{2}}\right]_{-u^{6}}^{u^{6}}$ (36)
Thus,
$\left|H_{u}(s)\right|$ $\leq\frac{1}{(\log u)^{2}}\int_{-u^{6}}^{u^{6}}\left|g^{\prime\prime}_{u}(y)\right|dy+O\left(\frac{1}{u^{12-3\delta-6\varepsilon}\log u}\right)$ (37)
and
$\int_{-u^{6}}^{u^{6}}\left|g^{\prime\prime}_{u}(y)\right|dy$ $=\int_{-u^{6}}^{u^{6}}\left|\left[1-\phi_{u}(y)\right]g^{\prime\prime}(y)-2\phi^{\prime}_{u}(y)g^{\prime}(y)-\phi^{\prime\prime}_{u}(y)g(y)\right|dy$ (38)
$\leq\int_{-u^{6}}^{u^{6}}\left|\left[1-\phi_{u}(y)\right]g^{\prime\prime}(y)\right|+2\left|\phi^{\prime}_{u}(y)g^{\prime}(y)\right|+\left|\phi^{\prime\prime}_{u}(y)g(y)\right|dy$
For large $|y|$, we have,
$\left|\frac{1}{(1-\sigma-\delta+iy)(\sigma-\delta+iy)}\right|=O\left(\frac{1}{y^{2}}\right)$ (39)
and from the convexity property of the zeta function [titchmarsh1986], for $|y|\gg|t|$,
$\left|\zeta(1-\delta+i(t+y)\right|$ $=O\left(|y|^{\delta/2+\varepsilon}\right),$ (40)
$\left|\zeta^{(m)}(1-\delta+i(t+y)\right|$ $=O\left(|y|^{\delta/2+\varepsilon}(\log|y|)^{m}\right),\qquad m\in\mathbb{N}$
This yields,
$g,g^{\prime},g^{\prime\prime}\in L^{1}(\mathbb{R})$ (41)
where $L^{1}(\mathbb{R})$ denotes the space of absolutely integrable functions on the real line.
Therefore,
$\int_{-u^{6}}^{u^{6}}\left|\left[1-\phi_{u}(y)\right]g^{\prime\prime}(y)\right|dy\leq\int_{r(u)\leq|y|\leq u^{6}}\left|g^{\prime\prime}(y)\right|dy=O(1)$ (42)
Since $\max|\phi^{\prime}_{u}(y)|=O(1/w(u))$ and $\max|\phi^{\prime\prime}_{u}(y)|=O(1/(w(u))^{2})$, then we have,
$\int_{-u^{6}}^{u^{6}}\left|\phi^{\prime}_{u}(y)g^{\prime}(y)\right|dy\leq\max|\phi^{\prime}_{u}(y)|\int_{r(u)\leq|y|\leq r(u)+w(u)}\left|g^{\prime}(y)\right|dy\leq O\left(\frac{1}{w(u)}\right)O\left(w(u)\right)=O(1),$ (43)
$\int_{-u^{6}}^{u^{6}}\left|\phi^{\prime\prime}_{u}(y)g(y)\right|dy\leq\max|\phi^{\prime\prime}_{u}(y)|\int_{r(u)\leq|y|\leq r(u)+w(u)}\left|g(y)\right|dy\leq O\left(\frac{1}{(w(u))^{2}}\right)O\left(w(u)\right)=O\left(\frac{1}{w(u)}\right)$ (44)
Thus,
$\int_{-u^{6}}^{u^{6}}\left|g^{\prime\prime}_{u}(y)\right|dy=O\left(\frac{1}{w(u)}\right)=O\left(\sqrt{\log u}\right)$ (45)
Substituting by equation (45) in equation (37) leads to,
$|H_{u}(s)|=O\left(\frac{1}{(\log u)^{3/2}}\right)$ (46)
For each $\zeta(\sigma+it)=\zeta(1-\sigma+it)=0$, we select $0<\delta<\min(\sigma,1-\sigma)$ such that $\zeta(1-\delta+it)\neq 0$ (this can also be achieved by choosing $\delta$ such that $(1-\delta+it)$ at the same $t$ lies in the zero-free region of the zeta function guaranteed by [Mossinghoff2015], [Gourdon2004]), and accordingly $g(0)\neq 0$, and $A\neq k\pi,k\in\mathbb{N}$, so, $\sin A\neq 0$. Consequently, from equations (34) and (46), we have
$|I_{u}(s)|\asymp\left|\frac{2u^{1-\delta+it}g(0)\sin A}{\log u}\right|=\frac{2u^{1-\delta}\left|g(0)\sin A\right|}{\log u}$ (47)
Therefore, $\lim_{u\to\infty}|I_{u}(s)|$ diverges leading to a contradiction with equation (28). This implies that the initial assumption that $\zeta(s)=\zeta(1-\bar{s})=0$ for $\sigma\neq 1/2$ cannot hold. That is, all the non-trivial zeros of the zeta function must lie on the critical line, $\Re(s)=1/2$, concluding the proof of the Riemann hypothesis. ∎